# The Galactic population of canonical pulsars II

### v1

@@ Abstract
Context. Pulsars are highly magnetized rotating neutron stars, emitting in a broad electromagnetic energy range. These objects have been discovered more than 50 years ago and are astrophysical laboratories for studying physics at extreme conditions. Reproducing the observed population of pulsars refines our understanding of their formation, evolution as well as radiation processes and geometry.
Aims. In this paper, we improve our previous pulsar population synthesis by studying the impact of the Galactic gravitational potential and of the death line putting emphasis on the $\gamma$-ray pulsar population. In order to elucidate the necessity of a death line, refined initial distributions for spin periods and radial positions at birth were implemented, elevating the sophistication of our previous simulations to the most recent state-of-the-art.
Methods. The individual motion of pulsars in the Galactic potential is evolved by a fourth order symplectic integration scheme. Our pulsar population synthesis takes into account the secular evolution of a force-free magnetosphere and the magnetic field decay. Each pulsar is evolved from its birth up to the present time. The radio and $\gamma$-ray emissions are modeled respectively by the polar cap geometry and the striped wind model.
Results. When simulating a cohort of ten million pulsars akin to previous endeavors, the integration of a death line proves essential for a better agreement with observational trends. Yet, discrepancies emerges when modeling a reduced ensemble of only one million pulsars, endowed with a higher birth rate. This divergence yields a remarkably refined $P-\dot{P}$ diagram, conspicuously with or without any delineating death line. Intriguingly, this observation suggests an overestimation of the ages of detected pulsars in existent data sets, thereby challenging the necessity of death lines in pulsar population synthesis. Validation through Kolmogorov-Smirnov testing underlines the statistical congruence between the spin period and spin period derivative distributions of both observations and simulations. Moreover simulations with a higher $\gamma$-ray instrumental sensitivity show that the GeV excess in the Galactic centre could be of pulsar origin.
@@ 1 Introduction
Pulsars, discovered by Jocelyn Bell Burnell in 1967 (Hewish et al., 1968) are rotating neutron stars born in a core-collapsed supernova. Pulsars are highly magnetized and surrounded by a plasma-filled magnetosphere emitting regular pulses of radiation at their spin frequency. Due to the magneto-dipole losses, they lose energy as their spin period increases. Thanks to radio surveys, performed for instance by the Parkes and Arecibo radio-telescopes, nearly 3000 pulsars have been discovered (Manchester et al., 2005) (see Table 1, with the data on the canonical pulsars). Although pulsars were first observed in radio, it was found later that they are also bright in X-rays, optical and gamma-rays. The Large Area Telescope (LAT) on board of the Fermi satellite, has discovered dozens of radio-quiet gamma-ray pulsars as well as millisecond pulsars (MSPs). Since its launch, the LAT has detected about 300 pulsars (Abdo et al., 2013; Smith et al., 2019, 2023). The LAT has unveiled a new perspective, expanding our understanding about pulsars by presenting a growing collection of neutron stars that are detectable through their high energy emission. Moreover, the distinct emission process in high-energy and its unique beaming properties offer an alternative viewpoint on the global pulsar population.
@@
Table 1: Number of known pulsars with spindown luminosity $\dot{E}$ above and below $10^{28}$ W, and above $10^{31}$ W. The quantities $N_{r}$, $N_{g}$, and $N_{rg}$ are the number of radio-only, gamma only, and radio-loud gamma-ray pulsars, respectively. It should be noted that we excluded the binary pulsars as well as pulsars with $P<20$ ms. The data have been taken from the ATNF catalogue and from https://confluence.slac.stanford.edu/display/ GLAMCOG/Public+List+of+LAT-Detected+Gamma-Ray+Pulsars
log($\dot{E}$) (W) $N_{tot}$ $N_{r}$ $N_{g}$ $N_{rg}$
¿31 2 0 0 2
¿28 197 101 35 61
Total 2665 2553 63 84
@@
With the advent of future surveys such as the Square Kilometre Array (SKA  [^1] https://www.skao.int/en/science-users/118/ska-telescope-specifications) or the Cherenkov Telescope Array (CTA  [^2] https://www.cta-observatory.org/about/how-ctao-works/), much more pulsars are expected to be detected. Those instruments are scheduled to start collecting data by the end of the decade. SKA is going to be 50 times more sensitive and is predicted to survey the sky 10 000 times faster than any existing imaging radio telescope (McMullin et al., 2020). CTA is aimed at detecting gamma rays in the energy range from a few tens of GeV up to hundreds of TeV (Hofmann & Zanin, 2023) while Fermi/LAT operates in the energy range from 20 MeV to several hundred GeV  [^3] https://fermi.gsfc.nasa.gov/science/instruments/table1-1.html.
@@
Pulsar population synthesis (PPS) is a powerful tool aiming at predicting the discovery rate of new pulsars as well as better understanding their emission processes. PPS studies also constrains the overall properties of pulsars. Pulsars are generated from their birth and evolved up to the present. Once the set of pulsars has been generated we check whether they fulfill some prescribed detection criteria. The detectability of an individual pulsar in a given energy band (radio or $\gamma$) depends on whether the associated emission beam intersects our line of sight and on the sensitivity of the corresponding instrument.
@@
Most of PPS studies assume that neutron stars are rotating in vacuum (Faucher-Giguère & Kaspi, 2006; Popov et al., 2010; Johnston et al., 2020), or take only radio or gamma-ray emission into account (Watters & Romani, 2011; Gullon et al., 2014), or assume a constant magnetic field (Gonthier et al., 2002; Faucher-Giguère & Kaspi, 2006; Johnston & Karastergiou, 2017). For the first time in a PPS study, Dirson et al. (2022) have modeled the population of both the gamma-ray and radio pulsars by taking into account the force-free magnetosphere model in conjunction with the magnetic field decay. The neutron star radiation is produced by relativistic charged particles flowing within their magnetosphere. Therefore it is necessary to take the plasma back reaction into account for the pulsar period and dipole magnetic inclination angle evolution.Whereas the gamma-ray emission site from pulsars is still under debate, different locations have been suggested. These are for instance the polar cap region (Sturrock, 1971; Ruderman & Sutherland, 1975; Daugherty & Harding, 1982, 1996) at low altitudes, the slot gap along the last open magnetic field lines (Arons, 1983; Muslimov & Harding, 2004; Harding et al., 2008; Harding & Muslimov, 2011), or the outer gaps at high altitudes in the outer magnetosphere (Cheng et al., 1986; Hirotani, 2008; Takata et al., 2011). The striped wind model, which has for emission site the pulsar wind outside the light cylinder, (Pétri, 2009, 2011) is used for the first time by Dirson et al. (2022) in a PPS study. The radio emission is described as usual by the polar cap model.
In this work, and the work of Dirson et al. (2022) only the canonical pulsar population in our Galaxy is reproduced. The canonical pulsar population is defined by isolated pulsars which are not magnetars, and with a spin period greater than 20 ms.
@@
While the study yielded satisfactory results when simulating a population closely resembling the observed pulsar population in the $P-\dot{P}$ diagram, there was room for improvement regarding the pulsar motion. Specifically, the study is limited by the assumption that pulsars move in straight lines at a constant speed from their birth to the present time. A noticeable discrepancy arises when comparing the observed position of pulsars to the simulation expectations, revealing a significant disparity in their spatial distributions.
Since the results obtained by Dirson et al. (2022) did not require the modeling of a death line, it was not included in their PPS. The death line separates in the $P-\dot{P}$ diagram the neutron stars that generate particles through pair cascading from those which cannot radiate anymore and are hence called dead pulsars. This contrasts with the PPS performed by Graber et al. (2023) who did not require a death line. Therefore in this paper we examine the necessity of a death line to explain the observed canonical pulsars population.
@@
This paper improves the PPS of Dirson et al. (2022), questions the necessity of a death line in PPS work and discusses in depth the $\gamma$-ray pulsar population. The paper is organised as follows. In Sect. 2 we detail our model, recalling the processes for generation and evolution of the pulsars in the PPS, a description of the Galactic potential, the death line and the multi-wavelength detection criteria. Then the results are shown in Sect. 3. The paper ends with a discussion in Sect. 4 and a summary in Sect. 5.
@@ 2 Description of the PPS model
Our PPS study is based on a self-consistent state-of-the-art evolution of the pulsar geometry and proper motion within the Milky Way. In this section we detail the model by exposing first the generation of individual pulsars, their initial position, magnetic field and spin period, then how these quantities should evolve in time, giving explicit expressions for the Galactic potential used in our simulations, before introducing the death line and finally the radio and $\gamma$-ray detection criteria.
@@ 2.1 Generation of pulsars
We generate a population of pulsars whose ages are generated in ascending order, every X years for example if the birth rate is 1/X yr-1 between 0 yr and the age of the Milky Way at maximum. The age of the Milky Way is almost equal to the age of the Universe, which is of about $t_{H}\approx 13.8\times 10^{9}$ yr, therefore, a total number of $t_{H}/X$ pulsars should be simulated. However as we will see, we chose to emulate $10^{7}$ pulsars in a first run and $10^{6}$ in another run. This choice was made for two reasons : firstly, the computation time is significantly decreased while pulsars older than $10^{8}$ yr are mostly not detected, secondly, the simulation that emulates $10^{7}$ pulsars allows to emulate older pulsars and require a death line while the other works well without. This approach was also chosen because it allows to check whether our birth rate is relevant by comparing our results from the simulation to the observations. It is different from other works like Johnston & Karastergiou (2017) and Gullon et al. (2014), where they stop the simulation when the number of pulsars detected in the simulation are the same as the number of observed pulsars. The value chosen for the birth rate, which is 1/41 yr-1, is taken as best educated guess from the interval [1/150,1/33] yr-1 because these studies (Faucher-Giguère & Kaspi, 2006; Gullon et al., 2014; Johnston & Karastergiou, 2017) evaluated the birth rate of pulsars in this interval.
To describe the position of the pulsars in the Galaxy, we use the right-handed Galacto-centric coordinate system $(x,y,z)$ with the Galactic centre at its origin, $y$ increasing in the disc plane towards the location of the Sun, and $z$ increasing towards the direction of the north Galactic pole. The initial spatial distribution of the pulsars is given by Paczynski (1990). He found two distributions, one describing the radial spread and one the spread in altitude, however only his distribution in altitude is used in our work because the radial distribution in the galactic plane is better depict in Ahlers et al. (2016) by the Milky Way’s pulsar surface density defined as
@@
where $R$ is the axial distance from the $z$-axis, and $z$ is the distance from the Galactic disc. The numerical values for the constants are, $A=37.6$ kpc-2, $a=1.93$, $b=5.06$, $R_{1}=0.55$ kpc, $h_{c}=180$ pc and $R_{\odot}=8.5$ kpc and in agreement with the distribution of young massive stars in our galaxy (Li et al., 2019). The purpose of choosing equation (1a) to describe the radial spread is to put neutron stars birth positions within the spiral arms of the Galaxy. More precisely the Galactic spiral structure contains four arms with a logarithmic shape function allowing to obtain the azimuthal coordinate $\phi$ as a function of the distance from the Galactic center
$\phi(R)=k\ln\left(\frac{R}{r_{0}}\right)+\phi_{0}$ (2)
@@
Table 2: Parameters of the Milky Way Spiral Arm structure.
@@
Next, we need to describe the individual properties of each pulsar, namely, its inclination angle $\alpha$, which is the angle between its rotation axis and its magnetic axis, its initial spin period and initial magnetic field at birth. The inclination angle $\alpha$ is assumed to follow an isotropic distribution generated from a uniform distribution $U\in[0,1]$ and given by $\alpha=\arccos(2\,U-1)$. In our work, the initial spin period and magnetic field follow both a log-normal distribution, as suggested from the results of a study of 56 young neutron stars by Igoshev et al. (2022) while other works simply use a Gaussian distribution (Faucher-Giguère & Kaspi, 2006; Gullon et al., 2014; Johnston et al., 2020). Explicitly, the probabilities to born with an initial magnetic field $B_{0}$ and period $P_{0}$ are given by
@@
Compared to the study of Dirson et al. (2022) we introduce a novelty, namely the alignment between the kick velocities at birth and the rotation axis of the pulsar, as suggested by Rankin (2007). Moreover according to Hobbs et al. (2005) a Maxwellian distribution best replicates the observations, with a standard deviation $\sigma_{\rm v}=265$ km/s. The mean velocity of the distribution is related to the standard deviation by $\bar{v}=\sigma_{\rm v}\sqrt{{8}/{\pi}}$. The velocity vector is then distributed along the unit vector of the rotation axis generated as detailed in the subsection 2.5.

### v2

@@ Abstract
Context. Pulsars are highly magnetized rotating neutron stars, emitting in a broad electromagnetic energy range. These objects have been discovered more than 55 years ago and are astrophysical laboratories for studying physics at extreme conditions. Reproducing the observed pulsars population refines our understanding of their formation and evolution scenarios as well as their radiation processes and geometry.
Aims. In this paper, we improve our previous population synthesis by focusing on both the radio and $\gamma$-ray pulsar populations, investigating the impact of the Galactic gravitational potential and of the radio emission death line. In order to elucidate the necessity of a death line, refined initial distributions of spin period and spacial position at birth were implemented, elevating the sophistication of our simulations to the most recent state-of-the-art.
Methods. The motion of each individual pulsar is tracked in the Galactic potential by a fourth order symplectic integration scheme. Our pulsar population synthesis takes into account the secular evolution of the force-free magnetosphere and magnetic field decay simultaneously and self-consistently. Each pulsar is evolved from its birth up to the present time. The radio and $\gamma$-ray emission locations are modelled respectively by the polar cap geometry and the striped wind model.
Results. By simulating ten million pulsars we found that including a death line better reproduces the observational trend. However, when simulating one million pulsars, we obtain an even more realistic $P-\dot{P}$ diagram, whether or not a death line is included. This suggests that the ages of the detected pulsars might be overestimated, therefore questioning the real need for a death line in pulsar population studies. Kolmogorov-Smirnov tests confirm the statistical similarity between the observed and simulated $P-\dot{P}$ diagram. Additionally, simulations with increased $\gamma$-ray telescope sensitivities hint to a significant contribution of $\gamma$-ray pulsars to the GeV excess in the Galactic centre.
@@ 1 Introduction
Pulsars, discovered by Jocelyn Bell Burnell in 1967 (Hewish et al., 1968) are rotating neutron stars born in a core-collapse supernova. They are highly magnetized and surrounded by a plasma-filled magnetosphere emitting regular pulses of radiation at their spin frequency. Due to the magneto-dipole losses, they lose rotational kinetic energy and their spin period increases. Extensive radio surveys, performed for instance by the Parkes and Arecibo radio-telescopes, so far 3700 pulsars have been discovered (Manchester et al. 2005, see Table 1 listing the canonical pulsars). Although pulsars were first observed in radio, they were later found to be also bright in X-rays, optical and $\gamma$-rays. The Large Area Telescope (LAT) on board of the Fermi satellite, has discovered dozens of radio-quiet $\gamma$-ray pulsars as well as millisecond pulsars (MSPs). Since its launch, the LAT has detected about 300 pulsars (Abdo et al., 2013; Smith et al., 2019, 2023) and unveiled a new perspective, expanding our understanding about pulsar emission mechanisms by increasing the sample of neutron stars detectable through their high energy emission. Moreover, the distinct radiation process involved in high-energy photon production and its unique beaming properties offer an alternative viewpoint on the global pulsar population.
@@
Table 1: Number of known canonical pulsars, with spindown luminosity $\dot{E}$ above and below $10^{28}$ W, and above $10^{31}$ W. The quantities $N_{\rm r}$, $N_{\rm g}$, and $N_{\rm rg}$ are the number of radio-only, gamma only, and radio-loud $\gamma$-ray pulsars, respectively. It should be noted that we excluded the binary pulsars, the pulsars in the Magellanic clouds and in globular clusters. The data have been taken from the ATNF catalogue from https://www.atnf.csiro.au/research/pulsar/psrcat/
log($\dot{E}$) (W) $N_{\rm tot}$ $N_{\rm r}$ $N_{\rm g}$ $N_{\rm rg}$
¿31 4 0 2 2
¿28 150 40 44 66
Total 2193 2020 80 93
@@
With the advent of future surveys such as the Square Kilometre Array (SKA  [^1] https://www.skao.int/en/science-users/118/ska-telescope-specifications) or the Cherenkov Telescope Array (CTA  [^2] https://www.cta-observatory.org/about/how-ctao-works/), much more pulsars will be detected in both these wavelengths. Those instruments are scheduled to start collecting data by the end of the decade. SKA is going to be 50 times more sensitive than current telescopes and will survey the sky 10,000 times faster than any existing imaging radio telescope (McMullin et al., 2020). CTA is aimed at detecting $\gamma$-rays in the energy range from a few tens of GeV up to hundreds of TeV (Hofmann & Zanin, 2023) while Fermi/LAT operates in the energy range from 20 MeV to several hundred GeV  [^3] https://fermi.gsfc.nasa.gov/science/instruments/table1-1.html and H.E.S.S operates in the energy range from 0.03 to 100 TeV  [^4] https://www.mpi-hd.mpg.de/HESS/pages/about/.
@@
Pulsar population synthesis (PPS) is a powerful tool to predict the discovery rate of new pulsars, to better understand their emission processes and to constrain their overall properties. In a PPS, pulsars are generated from their birth and evolved up to the present time. Once the sample of pulsars is generated, detection criteria are applied to check whether they can be detected by current and future telescopes. The detectability of an individual pulsar in a given energy band, radio or $\gamma$-ray, depends on whether the associated emission beam intersects our line of sight and on the sensitivity of the corresponding instrument.
@@
Most PPS studies assume that neutron stars are rotating in vacuum (Faucher-Giguère & Kaspi, 2006; Popov et al., 2010; Johnston et al., 2020) or take only radio or $\gamma$-ray emission into account separately (Watters & Romani, 2011; Gullon et al., 2014), or assume a constant magnetic field during the evolution process (Gonthier et al., 2002; Faucher-Giguère & Kaspi, 2006; Johnston & Karastergiou, 2017). Dirson et al. (2022) have modelled for the first time the population of both the $\gamma$-ray and radio pulsars, taking into account the state-of-the-art force-free magnetosphere model in conjunction with a prescription for the magnetic field decay. The neutron star radiation is produced by relativistic charged particles flowing within their magnetosphere. Therefore the plasma back reaction must be taken into account in the evolution of the pulsar period and magnetic inclination angle. Whereas the location of the $\gamma$-ray emission site is still under debate, different regions have been suggested. These are for instance the polar cap region (Sturrock, 1971; Ruderman & Sutherland, 1975; Daugherty & Harding, 1982, 1996) at low altitudes, the slot gap along the last open magnetic field lines (Arons, 1983; Muslimov & Harding, 2004; Harding et al., 2008; Harding & Muslimov, 2011), or the outer gaps at high altitudes within the light-cylinder, in the outer magnetosphere (Cheng et al., 1986; Hirotani, 2008; Takata et al., 2011). Emission is also possible outside the light cylinder, in the so-called striped wind (Pétri, 2009, 2011) and was used for the first time by Dirson et al. (2022) in a PPS study. The radio emission is described as usual by the polar cap model. In this work as in Dirson et al. (2022) only the canonical pulsar population in our Galaxy is reproduced. The canonical pulsar population is defined by isolated pulsars which are neither magnetars nor millisecond pulsars.
@@
While the study by Dirson et al. (2022) yielded satisfactory results when simulating a population closely resembling the observed pulsar population in the $P-\dot{P}$ diagram, there was room for improvement regarding the pulsar trajectories within the Milky Way. Specifically, Dirson et al. (2022) study was limited by the assumption that pulsars move in a ballistic motion depicted by straight lines at a constant speed from their birth to the present time. A noticeable discrepancy arises when comparing the observed position of pulsars to the predictions, revealing a significant disparity in both spatial distributions.
Dirson et al. (2022) did not require a death line to reproduce the $P-\dot{P}$diagram and was therefore not implemented in their PPS. However the death line segregates neutron stars generating particles through pair cascading from those not radiating any more because of the quenching of this pair avalanche, and hence they are called dead pulsars. The effect of the death line assumption on the PPS differs from the PPS performed by Graber et al. (2024) who did not consider a death line but instead relied only on as special radio luminosity law to eliminate some pulsars in the $P-\dot{P}$ diagram.
@@
In this paper we further improve the PPS of Dirson et al. (2022) by examining the necessity of a death line to explain the observed canonical pulsars population and discuss in depth the $\gamma$-ray pulsar population. Moreover, our study helps to better understand the impact of changing the parameters of our model as well as their physically meaningful ranges. The paper is organised as follows. In Sect. 2 we detail our PPS model, recalling the processes for generation and evolution of the pulsar sample, the description of the Galactic potential and of the death line, discussing our adopted multi-wavelength detection criteria. The results are shown in Sect. 3. The results are discussed in Sect. 4 and Sect. 5 covers a summary of the results.
@@ 2 Description of the PPS model
Our PPS study is based on a self-consistent state-of-the-art evolution of the neutron star geometry of radio and $\gamma$-ray emission, magnetic field and proper motion within the Milky Way. In this section we detail our model by exposing, first the generation of individual pulsars, their initial position, magnetic field and spin period, second how these quantities evolve in time. We also give explicit expressions for the Galactic potential used in our simulations and introduce the death line before discussing the radio and $\gamma$-ray detection criteria. We stress that the model parameters used in this section have been carefully chosen by comparing simulated data with actual observations, varying the unknown input parameters in sensible physical ranges (see appendix C for more details).
@@ 2.1 Generation of pulsars
We generate a population of pulsars whose real ages are selected in ascending order, for example every X years if the birth rate is 1/X yr-1, with their age being between 0 yr and the age of the Milky Way at maximum. The age of the Milky Way does not significantly differ from the age of the Universe estimated about $t_{H}\approx 13.8\times 10^{9}$ yr, therefore, a total number of $t_{H}/X$ pulsars should be simulated. The best value for this birth rate was found to be 1/41 yr-1, see appendix C. Our value is consistent with earlier investigations by Faucher-Giguère & Kaspi (2006); Gullon et al. (2014); Johnston & Karastergiou (2017) where they evaluated the birth rate of pulsars to be in the range [1/150,1/33] yr-1. With $X=41$ yr we should generate more than $10^{7}$ pulsars, however we chose to emulate only $10^{7}$ pulsars in a first run and $10^{6}$ in a second run for the following two reasons. Firstly, the computation time is significantly decreased, secondly pulsars older than $10^{7}$ yr are mostly not detected. This approach was also chosen because it allows us to verify whether our birth rate is realistic, by comparing our results from the simulation to the observations. Our approach is different from previous works like Gullon et al. (2014) and Johnston & Karastergiou (2017), where they stop the simulation whenever their detected number of pulsars equals the number of observed pulsars.
To describe the position of the pulsars in the Galaxy, we use the right-handed Galacto-centric coordinate system $(x,y,z)$ with the Galactic centre at its origin, $y$ increasing in the disc plane towards the location of the Sun, and $z$ increasing towards the direction of the north Galactic pole. The initial spatial distribution of the pulsars is given by Paczynski (1990). He found two distributions, one describing the radial spread and one the spread in altitude, however only his distribution in altitude is used in our work because the radial distribution in the galactic plane is better depicted in Yusifov & Küçük (2004) by the Milky Way’s pulsar surface density defined as
@@
where $R$ is the axial distance from the $z$-axis, and $z$ is the distance from the Galactic disc. The numerical values for the constants are, $A=37.6$ kpc-2, $a=1.64$, $b=4.0$, $R_{1}=0.55$ kpc, $h_{c}=180$ pc and $R_{\odot}=8.5$ kpc and in agreement with the distribution of young massive stars in our galaxy (Li et al., 2019). The purpose of choosing equation (1a) to describe the radial spread is to put neutron stars birth positions within the spiral arms of the Galaxy. More precisely the Galactic spiral structure contains four arms with a logarithmic shape function allowing to obtain the azimuthal coordinate $\phi$ as a function of the distance from the Galactic center
$\phi(R)=k\ln\left(\frac{R}{r_{0}}\right)+\phi_{0}.$ (2)
@@
Table 2: Parameters of the Milky Way Spiral Arm structure, adapted from Yao et al. (2017).
@@
Next, we need to describe the individual properties of each pulsar, namely, its inclination angle $\alpha$, which is the angle between its rotation axis and its magnetic axis, its initial spin period $P_{0}$ and initial magnetic field $B_{0}$ at birth. The inclination angle $\alpha$ is assumed to follow an isotropic distribution generated from a uniform distribution $U\in[0,1]$ and given by $\alpha=\arccos(2\,U-1)$. In our work, the initial spin period and magnetic field follow both a log-normal distribution, as suggested from the results of a study of 56 young neutron stars by Igoshev et al. (2022) while other works simply use a Gaussian distribution for the spin period at birth (Faucher-Giguère & Kaspi, 2006; Gullon et al., 2014; Johnston et al., 2020) while for the magnetic field the prescription is the same as ours. Explicitly, the probability distributions for an initial magnetic field $B_{0}$ and an initial period $P_{0}$ are given by
@@
According to Hobbs et al. (2005) a Maxwellian distribution for the kick velocity at birth best replicates the observations and is given by
$p(v)=\sqrt{\frac{2}{\pi}}\frac{v^{2}}{\sigma_{v}^{3}}\exp{\left(-\frac{v^{2}}{2\sigma_{v}^{2}}\right)}$ (5)
The mean velocity of the distribution is related to the standard deviation by $\bar{v}=\sigma_{\rm v}\sqrt{{8}/{\pi}}$ with a standard deviation $\sigma_{\rm v}=265$ km/s for pulsars. The velocity value drawn from the Maxwellian distribution is then distributed along the unit vector of the rotation axis which is generated, as detailed in subsection 2.5. Moreover compared to the study of Dirson et al. (2022) we introduce a novelty, namely the alignment between the kick velocities at birth and the rotation axis of the pulsar, as suggested by Rankin (2007).